Home‎ > ‎Reports‎ > ‎Crew102‎ > ‎

day14


Crew 102 Log Book for April 8, 2011

Photos of the day





Commander's Check-In Report

Franco Carbognani Reporting

Crew Physical Status: All the Crew is doing fine.

Time Departed/Returned from EVA: 12:00 - 14:10, 16:30-18:30

Brief Narrative of Field Mission Results:

Collected cyanobacteria samples

Collected halophiles samples 

EVA Data/Interpretations:

Cyanobacteria and halophiles samples will be further analyzed at AMES laboratory 

Engineering/Hab Maintenance:

Downstairs furnace still off but wall mounted heater backup is fine.

Plans for Tomorrow:

Cleaning, packing all scientific equipment, getting ready for traveling back to Grand Junction.

Handover to Crew 103

Support Requested:

Don Lusco for:

1) Connect back the Hab gray water piping to the gray water tank

2) Fix downstairs furnace

Miscellaneous:


Commander's Report

Franco Carbognani Reporting

Today has been busy with the last 2 EVAs and with summary report writing.

This has been for all of us a wonderful experience and is with some sadness that we are getting prepared

for the next MDRS Crew handover.

But the compelling question on future astronauts feed and oxygen is still pending.

Our spirulina cultures did not jet gave a definite reply

Our Crew biologist will keep analyzing those cultures once back on her home laboratory.

So, stay tuned for the final results.

Also, many other studies we have started here at MDRS will sparkle follow-up projects and initiatives.

So this is the beginning and not the end.

Long life to Crew 102!


Engineering Report

Franco Carbognani Reporting

Generator/Electricity:

Kitty (Diesel Generator) (Used? Yes/No): Yes

Honey (Gas Generator) (Used? Yes/No): No

Battery Charge Level:


Notes/Comments (include how many generator runs, approximately when, and any times you needed to turn the generator on/off):


Fuel Status:

Diesel (Full/0.75/0.5/0.25/Empty): 0.55

Propane Tank (Full/0.75/0.5/0.25/Empty): 0.46

Gasoline Tank (Full/0.75/0.5/0.25/Empty): 0.54

Notes/Comments:


Oil:

Kitty Oil Quantity (# of Quarts in storage): 4/4

Honey Oil Quantity (# of Quarts in storage): 0/4, to be refilled

ATV Oil Quantity (# of Quarts in storage): all 4/4

Car Oil Quantity (# of Quarts in storage): 3/4


Water Status:

Outside Potable Water Tank Level (inches from bottom): 10

Trailer Potable Water Tank Level (inches from bottom): full

Hab Potable Water Tank Level (inches from bottom): 20

Potable Water Meter Reading: 44830,8

In to GreenHab Meter Reading:

Notes/Comments:


GreenHab:

Condition of plants in Tank 1:

Condition of plants in Tank 2:

Condition of plants in Tank 3:

Greenhab Notes/Comments:

Gray Water System: The Trickle Reservoirs are filled with gray water

Pumps running fine and water circulating. Acquatic tanks not refilled for lack of gray water.


Observatory:

Telescope Used? (Yes/No):No

Observatory Notes/Comments:


Transportation:

Hab Car (NOT RENTAL)

Used (yes/no): no

Oil Checked (yes/no):no

Tire Status: good

Odometer at end of day: 123903.8

Notes/Comments on Hab Car:


ATV 1 (Opportunity):

Used: yes

Oil Checked: no

Fuel Consumed: 1/2 liter

Tire Status: inflated

Comments:


ATV 2 (Spirit):

Used: yes

Oil Checked: no

Fuel Consumed: 1/2 liter

Tire Status: inflated

Comments:


ATV 3 (Viking I):

Used: yes

Oil Checked: no

Fuel Consumed: 1/2 liter

Tire Status: good

Comments:


ATV 4 (Viking II):

Used: no

Oil Checked: no

Fuel Consumed: none

Tire Status: to be inflated

Comments: this ATV does not start


Heating and Ventilation:

Thermostat setting upstairs (Farenheit): 65

Actual temperature upstairs (Farenheit): 77

Thermostat setting downstairs (Farenheit): N/a

Actual temperature downstairs (Farenheit): 67

Heating and Ventilation Notes/Comments:

Downstairs furnace has stopped most probably for a fan not working properly.

We are using as a backup the wall mounted heater which is working fine.


Computers/Networking Infrastructure:

Notes/comments on internet/computers?:


General Engineering Notes/Comments:


Health and Safety Report

Dr. John E. Deaton Reporting

END OF SIMULATION REPORT

Crew #102 were resident at the MDRS from March 26 to April 9, sim ending April 8. During that time the following events associated with Health and Safety of the crew occurred—

• Crew briefed on safety features within the Hab and the location of medical supplies

• Safety inspections were conducted weekly

• HSO submitted regular HSO updates to Mission Support

• Crew complained of low grade intestinal issues; they did not resolve during our 2 week period, but they were manageable

• HSO experienced infection in the eyes, inflammation in the throat and ears, and moderate cough; a medical telecom was conducted with the Flight Surgeon who recommended the HSO break sim and drive to Bicknell to receive treatment (which he did, accompanied by a fellow crew member)

• Treatment was administered at Bicknell (antibiotics and cough medicine) and HSO is continuing to recover, and is nearly back to 100%

• French film crew member suffered what we think was heat exhaustion; she was treated and next day appeared to be fully recovered

• No other significant events occurred during this time period

Recommendations:

• Include antibiotic eye drops in medical supplies

• Confirm calibration of thermometer; appears not to be accurate; perhaps replace

• Repair/redesign toilet facility so crew members do not have to carry buckets of gray water up and down stairs; this is an accident waiting to happen!

• Rewiring of power strip; as it is now placed it is in direct traffic of crew walking on the second deck; several people have already tripped on this

Note from HSO: Telecom with Flight Surgeon was very successful, and the rapid response rate during this process was commendable!


EVA 16 Report

Kavya k.Manyapu Reporting

Time of Departure:  12.00 pm

Time of Arrival: 14:10 pm

Crew Members: L.Vimercati, K.Manyapu, A.Kapoglou

Site Location:  UTM 12S0520073 4251636

Altitude: 4536 feet

Total Distance: 6 miles

Transit Mode: ETV

Objectives: The EVA was the last expedition in search of cyanobacteria

Results:  The crew has been successful in marking the locations where they found cyanobacteria in abundance on the Mars analogue. Samples have been collected for further analysis beyond the rotation.

Lessons Learned:

It is recommended not to perform an EVA, especially on ETVs when the temperatures are low and the weather is very windy.


EVA 17 Report

Yuval Brodsky Reporting

Time: 1630 - 1830

Crew Members: Y. Brodsky, F. Carbognani, L. Vimercati

Site Location: UTM 12S 517000 4253000

Transit Mode: ATVs

Objectives: Halophile sample collection

Results: Numerous halophile samples successfully collected.

Lessons Learned: N/A


Journalist Report

Kavya Manyapu Reporting

And the day has finally arrived. Crew 102 has successfully completed their rotation on the Red planet and has accomplished several goals. Human Exploration of Mars definitely poses a great challenge. The first few settlements on Mars will require astronauts to be a well rounded people, don several hats-being an engineer, biologist, geologist, chef,  janitor, to quote a few.  MDRS is a great opportunity for researchers, scientists and astronaut hopefuls to participate, experience, learn and understand the operational constraints and what it is like living on Mars.

The ambitious crew has been working until the end of the day, trying to squeeze in couple of EVAs, run biology experiments even on the final day while packing and cleaning the Hab for the next crew’s arrival. The first EVA of the day was a final expedition in search of cyanobacteria. The crew for this expedition was all female marsonauts of the rotation, making it an interesting EVA. The EVA was commanded by L.Vimercati and accompanied by K.Manyapu and A.Kapoglou on the ETVs. The final EVA of the crew was in search of halophile samples commanded by Y.Brodsky , accompanied by L.Vimercati and F.Carboganani. 

Although an extremely diverse group and a very international crew, the crew dynamics have been exceptional during this mission. Each crew member not only was brilliant in their own area of expertise but also helped in various other areas during EVAs and experiments. Experiments in the areas of geology, biology, physiology, paleontology, material science, human factors and engineering were conducted by Crew 102 during their rotation. Along with 3 outreach experiment s for Space Florida, the crew has also coordinated with the French TV channel “ARTE” in filming a documentary, promoting Mars Society and MDRS which they hope will inspire many young children, students and young professionals to participate and promote human space exploration. The crew would like to spend more time on MDRS and they wish the simulation could have been longer in order to obtain more experience; more importantly to be able to conduct long term experiments to obtain viable data. The crew would like to surely team up once again and come back to MDRS.  

It is now time to see how the crew will re-adapt to civilization on earth and once again feel earthling.  

WHERE THERE IS WILL, THERE IS A WAY!